OpenStudy (anonymous):

Write 8/74 in simplest form.? A. 8/37 B.4/74 C.4/37 D.8/74

OpenStudy (anonymous):

??

OpenStudy (diyadiya):

\[\huge \frac{8}{74}= \frac{\frac{8}{2}}{ \frac{74}{2}}\]

OpenStudy (diyadiya):

8 and 74 are both divisible by 2 So we divide both numbers by 2 what do you get now ?

OpenStudy (anonymous):

4/37

OpenStudy (diyadiya):

Good :D

OpenStudy (anonymous):

:) thanks

OpenStudy (diyadiya):

Anytime ^_^

OpenStudy (anonymous):

I also got 4/37 :D